The pressure required to prevent liquid from entering a plug of a finely divided solid is twice as great for a liquid of surface tension 50 mN m"1, which completely wets the solid, as it is for a liquid of surface tension 70 mN m"1, which has a finite contact angle with the solid. Calculate this contact angle
1
Expert's answer
2020-05-24T02:37:34-0400

We know that

Wetability is inversely proportional to the surface tension

So pressure required will also be incersly proportional to surface tension

Hence surface tension of liquid will be 25 mN - m

Contact angle ,"\cos\theta =\frac{25}{70}=0.3571"



"\theta =\cos^{-1}=0.3571=69\degree"